LOFAR Newsletters December 2015

Distributed on 15 December 2015

 

Array status:

----------------------

-       38 stations operational in the Netherlands: 24 core stations and 14 remote stations.

-       9 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609.

-       3 new international stations have been built in Poland (PL610, PL611, and PL612). The data connection has not been established yet, but is in progress.

-       The HBA fields of CS013 have been rotated back to their originally planned orientation. Following the earlier replacement of the LCU and extensive testing of the new RSP driver, the station has been declared fully operational again on Monday 9 December.

-       The storm season has started in the Netherlands. A few LBAs and HBAs suffered from strong winds and have been repaired.

-       Routine station maintenance has been put on hold because of the advanced autumn season. Activities will start again in the spring 2016.

-       The overview of non-operational antenna elements for LBA and HBA is available here.

-       Station calibration: during the last two months, calibration tables for modes 3, 5, and 7 were updated for almost all remote and international stations. A successful test was performed preparing tables for mode 6 in a subsample of stations. A data collection run for mode 6 on all stations has been performed on 14 December.

 

Observing System Status:

------------------------------------ 

-       The observing system was mostly stable during the period October - December 2015. However, failures of COBALT nodes, and the instability of processing and LTA (ingest) systems caused delays in the exploitation of the Cycle 4&5 observing programmes and in digesting data out of CEP2.

-       The low(er) ingest speeds experienced towards the Juelich LTA location when including the German stations in the observing setups have been solved through a reconfiguration at Juelich.

  

Software Development Status (H. Holties / J. Annyas):

------------------------------------------

-       A station-only software release (2.13) has been carried out on November 2nd. Most notable changes were improvements in the specification of attenuation settings and in the robustness and speed of clock switches, needed for supporting 160 MHz modes. The latter are now available to LOFAR operations.

-       An intermediate software release (2.14) in preparation for CEP4 commissioning took place on December 7th.

-       The software team is currently working on the implementation of a dynamic scheduler and required components to allow utilization of the CEP4 job scheduler.

-       Work is underway to connect the new Polish stations to the LOFAR network and to set up a new LOFAR LTA site hosted at the Poznan compute centre.

-       The LTA ingest process is being adapted to allow third party data ingests, initially to be used by the EOR and Dragnet groups.

-       Jasper Annyas started as new head of the Software Support Group on November 16th.

 

 

 

CITT Status (T. J. Dijkema & E. Orru')

--------------------------------------------------------------------

-       The CITT2 is currently focusing on stress testing the Factor pipeline, the pipeline for HBA imaging using the facet calibration technique. In a busy week organized in week 50, about 10 fields passed the initial stages of this pipeline (and a major bug was fixed). The awimager2 and image domain gridder on the citt branch was updated and is currently being tested. The smart demixer has been commissioned, and delivers good results now. Other DPPP improvements include multithreading in gaincal.

 

-       Discussions on LBA calibration have started, and a small workshop  will take place in January (organized by the Surveys group).

 

-       The CITT has sparked a discussion on the LOFAR pipeline framework. Replacing this should improve RO's ability to adopt new pipelines. This has led to a project proposal for 2016, which is now submitted.

 

Observing Programmes and miscellaneous items:

--------------------------------------------------------------------

 

-       35 Cycle 5 proposals where submitted. 33 projects have been accepted. In total, 1598 observing and 1724 processing hours were allocated.  Long-term status was also awarded to some projects. An overview of the allocations is given here.

-       Cycle 5 observing programme: 17% complete. Schedule available here.

-       Cycle 4 observing programme: 93% complete. The rest will be observed with second priority during Cycle 5.

-       Cycle 3 observing programme: 91% complete - Unfinished projects will not continue during Cycle 5.

-       LOFAR user contributed scripts have been migrated to GitHub and are now available to the entire user community. You can find them in the repository. The repository is also documented in the latest release of the LOFAR Imaging cookbook.

-       A new version of the LOFAR Imaging Cookbook (v. 18.0) has been release on November 11 2015 and can be downloaded here.

-       The list of LOFAR papers (and associated statistics) is available here.

-       As of December 1 2015, the BBS database server named 'ldb001' is not used anymore. In users scripts, any reference to ldb001 should be replaced with the alias 'ldb', which points to the new database machine ldb003.

-       'The Third LOFAR Users meeting' and ‘'The 2016 LOFAR Community Science Workshop' will take place on 4-6 April 2016 in Zandvoort aan Zee, nearby Amsterdam. The registration will open in January, when the second announcement will be circulated.

 

CEP news:

----------------

-       The new CEP4 cluster is still undergoing acceptance tests and is expected to be delivered to LOFAR in January. It is planned to be ready for operations in March 2016.

 

-       CEP2:

  • Disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends.
  • During the past two months, a few locus nodes experienced issues and were temporarily taken out of production.
  • Cleanup of users data on CEP2 has started. Users have been requested to remove any data they own on CEP2 by the end of November 2015. Access will be restricted in the next few days.

 

-       CEP3:

  • Cluster info and schedule available here.
  • Monitoring of node usage is active. Statistics will be used to evaluate requests for extensions.

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

-       Next LSM's : 06/01, 20/01, 03/02, 17/02, 02/03

(note: all presentations given at the LSM are available here).

-       Next Stop days: 04/01, 02/02, 01/03

-       Next software roll outs: 25/01, 14/03

-       'The Third LOFAR Users meeting': 04/05

-       ‘'The 2016 LOFAR Community Science Workshop': 05-06/05

-       Next LOFAR bulletin: February 2016.

Design: Kuenst.    Development: Dripl.    © 2020 ASTRON